OpenStudy (anonymous):

Simplify [(5)/(4x^2y)]-[(y)/(14xz)]

OpenStudy (anonymous):

\[\frac{ 35z }{ 28x^2yz }-\frac{ 2xy^2 }{ 28x^2yz }\] is this right and then what do i do after?

OpenStudy (anonymous):

this is the original problem on equation \[\frac{ 5 }{ 4x^2y }-\frac{ y }{ 14xz }\]

OpenStudy (anonymous):

. 5 . . . . -y . . . . 35z - 2xy² ------ + ------ = -------------- 4x²y . . 14xz . . . .28x²yz

OpenStudy (anonymous):

ok so that cant be simplified anymore thanks